# Knowledge in multi-robot systems: an interplay of dynamics, computation and communication

### v1

@@ 4.3 The Surveillance Task
4.4 Approximate Point Gathering Task
In the approximate gathering task, we require robots to eventually meet up in the same space.
Instead of gathering at a precise point, robots need to ensure that their position and all other robots positions’ are contained in a specific open set, called the rendezvous area. To this extent, we need to formulate statements depending on the robots current position, thus extending the previously introduced epistemic language with positional atoms, representing the position where the robot currently sits. Furthermore, we now explicitly require the set of possible observed positions $\text{P}_{g(X)}$ (i.e., a subset of the domain of $g$ in Definition 2.1) to be finite over the exploration space $X$, as set in Remark 2.4. $\text{P}_{g(X)}$ is a finite subset of our space $\text{P}_{g(X)}\subseteq X$, such that $|\text{P}_{g(X)}|\in\mbox{$\mathbb{N}$}$, where $\vec{x}\in\text{P}_{X}^{\Pi}$ denotes a vector with a position for each robot. If $\text{P}_{g(X)}$ does not cover the whole space, then there are places where a robot simply cannot go.
We denote its positional atoms by $pos_{\mathbf{r}}(x)$ meaning that robot $\mathbf{r}$ is at some actual position $\mathbf{x}$ that is discretized to $x$. Note that, as $g$ is surjective, we might have $g_{\mathbf{r}}(\mathbf{x})=g_{\mathbf{r}}(\mathbf{x}^{\prime})$ even though $\mathbf{x}\neq\mathbf{x}^{\prime}$, meaning two different actual robot positions result in the same position for that robot. By construction, we assume that robots know their own position: $pos_{\mathbf{r}}(x)\leftrightarrow K_{\mathbf{r}}pos_{\mathbf{r}}(x)$. We accordingly extend the semantics by $\pi(pos_{\mathbf{r}}(x))$ to contain any point $(\sigma,[t])$ where robot $\mathbf{r}$ believes to be at the position $x$. We further equip the language with facts about the inclusion of points in open sets in the exploration space $\in_{x,U}$ denoting that and abbreviated as $x\in U$ collected in the set $\text{P}_{\tau}$. Note that, since the position and subset atoms are parameterized by points and open sets, we can form expressions connecting them. For example $pos_{\mathbf{r}}(x)\wedge(x\in U)$ holding, denotes that robot $\mathbf{r}$ believes to be position $x$ which is completely contained in area $U$.
Additionally we identify by $init(pos_{\mathbf{r}}(x))$ the initial canonical position of robot $\mathbf{r}$, meaning its starting position, and extend its semantics via $\pi(init(pos_{\mathbf{r}}(x))=\{(\sigma,t)\;|\;(\sigma,0)\vDash pos_{\mathbf{r}}(x)\}$. Lastly, we define a finite set of exploration atoms to denote possible rendezvous areas $\text{P}_{rz(X)}\in 2^{X}$ such that $\forall U,V\in\text{P}_{rz(X)}:U\cap V=\varnothing$ and $|\text{P}_{rz(X)}|\in\mbox{$\mathbb{N}$}$. The approximate point gathering is solved if all robots’ positions are contained in the same rendezvous area forever, assuming each robot starts in one rendezvous area. By the previous reasoning, the robots also know their initial starting position. Naturally, the robots need not do anything when all their initial canonical position are in the same rendezvous area. Otherwise they need to agree on one rendezvous area and move to there. They are allowed to change the rendezvous area finitely often, but eventually need to stick to one forever. We require it to be impossible for any robot to intersect two rendezvous areas at the same time, i.e., $pos_{r}(x)\wedge(x\in U)\implies(\bigwedge_{U\neq V\in\text{P}_{rz(X)}}\neg(x\in V))$. This resembles the notion of taking a unique choice, where we forbid a robot to choose two areas simultaneously.
The following grammar defines temporal epistemic logic formulas for approximate point gathering.
@@ 4.4 Approximate Point Gathering Task
where $p\in\text{P}_{\tau}\cup\text{P}_{g(X)}\cup\text{P}_{rz(X)}$, $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$. When obvious from the context, we will omit the subscript $X$.
The semantics of $\mathcal{L}_{gather}$ is analogous to the one of $\mathcal{L}_{exp}$ in Definition 4.3, with the appropriate extensions to the interpretation function as already set.
Definition 4.35 (Approximate Point Gathering).
Any robot $\mathbf{r}$ strives to intersect their canonical position together with all other robots’ positions in a rendezvous area $u\in\text{P}_{rz(X)}$. Thus we say a system $\mathcal{I}$ is consistent with approximate point gathering, iff the following two conditions hold.
  • (Eventual Gathering): There is a rendezvous area where all robots eventually meet forever
    $\mathcal{I}\vDash\bigvee_{U\in\text{P}_{rz(X)}}\bigwedge_{\mathbf{r}\in\Pi}\Diamond\Box\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U).$ (20)
  • (Starting Validity): If all robots are already contained in a rendezvous area initially, that is also the final rendezvous area
    $\begin{split}\mathcal{I}\vDash\bigwedge_{\begin{subarray}{c}U\in\text{P}_{rz(X)},\\ \vec{x}\in\text{P}_{g(X)}^{\Pi}\end{subarray}}\bigl{(}(\bigwedge_{r\in\Pi}init(&pos_{\mathbf{r}}(x_{\mathbf{r}})))\wedge(x_{\mathbf{r}}\in U)\bigr{)}\rightarrow\\ \bigl{(}\Diamond\Box\bigvee_{\vec{y}\in\text{P}_{g(X)}^{\Pi}}(\bigwedge_{\mathbf{r}\in\Pi}&pos_{\mathbf{r}}(y_{\mathbf{r}}))\wedge(y_{\mathbf{r}}\in U)\bigr{)}.\end{split}$ (21)
This strongly resembles the stabilizing agreement problem, epistemically characterized in [10]. It is defined by the following formulas, where $\mathcal{V}$ is the set of values to choose from, the atom $choose_{a}(v)$ denotes that agent $a$ chooses value $v$. $decide_{a}(v)$ is defined as $\Box choose_{a}(v)$, where $a,b$ are agents, i.e., robots with no dynamics as written in Example 2.3.
@@
We say that a stable choice system is consistent with stabilizing agreement if (Agreement) and (Validity) hold.
  • (Agreement) : There is a value such that every agent decides on that value
    $\mathcal{I}\vDash\bigvee_{v\in\mathcal{V}}\bigwedge_{a\in\Pi}\Diamond decide_{a}(v)$ (22)
  • (Validity) : An agent can only choose a known initial value of some agent
    $\mathcal{I}\vDash\bigwedge_{v\in\mathcal{V}}\bigl{(}choose_{a}(v)\rightarrow K_{a}\bigvee_{b\in\Pi}init_{b}(v)\bigr{)}$ (23)
Even though the two conditions look syntactically equivalent, we have to do some work to show that stabilizing consensus implies approximate gathering. First observe that covering a rendezvous area can be seen as that robot choosing that area. In that sense, rendezvous areas are the objects the robots choose from. As each robot initially sits in exactly one area, that area becomes the problem input. The last obstacle is reducing the strong validity, as its called in the distributed computing literature (see for example [11]), to the weak validity of approximate gathering.
@@
Consider a model $\mathcal{I}$ with an exploration space $X$, together with the language defined Definition 4.34, an arbitrary run $\sigma$ from $\mathcal{I}$. If $\mathcal{I}$ is consistent with stabilizing agreement, then it is also consistent with approximate point gathering.
Consider setting the set of possible choices $\mathcal{V}$ to be the set of possible rendezvous areas $\text{P}_{rz(X)}$, setting $choose_{\mathbf{r}}(U)\leftrightarrow\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U)$ and $decide_{\mathbf{r}}(U)=\Box choose_{\mathbf{r}}(U)$, and setting $init_{\mathbf{r}}(U)\leftrightarrow\bigvee_{x\in\text{P}_{g(X)}}init(pos_{\mathbf{r}}(x))\wedge(x\in U)$.
Proof 4.38.
We start by substituting for $choose$ and $init$ as defined in the equations Equation 22 and Equation 23. For (Agreement) we immediately arrive at (Eventual Gathering):
$\mathcal{I}\vDash\bigvee_{U\in\text{P}_{rz(X)}}\bigwedge_{\mathbf{r}\in\Pi}\Diamond\Box\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U).$ (24)
For (Starting Validity) we need to provide more reasoning, consider the following, where we already pre-pended the validity with a conjunction over all robots:
$\begin{split}\mathcal{I}\vDash\bigwedge_{\begin{subarray}{c}\mathbf{r}\in\Pi,\\ U\in\text{P}_{rz(X)}\end{subarray}}\bigl{(}(&\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U))\rightarrow\\ (K_{\mathbf{r}}&\bigvee_{\mathbf{r}^{\prime}\in\Pi}\bigvee_{x\in\text{P}_{g(X)}}init(pos_{\mathbf{r}^{\prime}}(x))\wedge(x\in U))\bigr{)}.\end{split}$ (25)
@@
for any $U\in\text{P}_{rz(X)}$ and any $\vec{x}\in\text{P}_{g(X)}^{\Pi}$. Choose any $U$, $\vec{x}$ which satisfies Equation 26, meaning a starting position vector that intersects over $U$. Note that there is only one such $U$ as initial rendezvous areas are unique.
Observe that, as robots know their own initial positions, Equation 26 implies the antecedent in Equation 25 for all robots $\mathbf{r}$:
@@
Therefore the consequent in Equation 25 is true for only one choice of $U$ (the same that satisfies Equation 26) and false for all other $V\in\text{P}_{rz(X)}$.
By Equation 24 there exists some $V$, $t$ such that
$\mathcal{I},(\sigma,t)\vDash\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in V)$ (29)
holds for all $t^{\prime}>t$ and for all robots $r$. We see that Equation 29 is precisely the antecedent of Equation 25, and, as Equation 25 is a validity, it is satisfied by the fixed choice for $V$, $t$. Since the consequent in Equation 25 is true only for $U$ as derived in Equation 28 (but Equation 25 is a validity), Equation 29 has to be false for any $W\neq U$. But by the previous reasoning we know that Equation 29 is true for at least one $V$ for all $t^{\prime}>t$, implying that $U=V$ for all $t^{\prime}>t$. We can now derive Equation 27 from Equation 29:
@@
meaning the robots gather at the initially intersected rendezvous area forever.
This has non-trivial consequences: the second depth broadcaster condition identified in [10] immediately becomes a sufficient condition for approximate point gathering. As an application, the luminous robot model with full visibility is equivalent to the iterated immediate snapshot model [1] in the distributed computing view and therefore satisfies said second depth broadcaster condition. By the above reasoning, luminous with full visibility is immediately sufficient for approximate gathering.
@@ 5 Conclusions and future work
We developed a framework that integrates seemlessly control, distributed systems and (epistemic and temporal) logical aspects of multi-robot systems, with applications to exploration, surveillance and gathering tasks. This has multiple implications and holds the promise of multiple future developments. One of the most striking implications is that the classical luminous robot model [12] with full visibility is very strong in a distributed computing sense: not surprisingly, there is a whole research field on establishing and maintaining communication in multi robot systems. Some, such as [7, 8] have leveraged epistemic planning and action models in order to keep a live communication. As we have identified a sufficient condition for approximate gathering, we can directly judge the correctness of an action model, and therefore the correctness of a protocol, by its ability to reach a given logical specification.
As a follow-up of this work, we are considering extending the logic framework provided here by enriching the language with deontic modalities expressing robots’ preferences (or strategies) during protocol execution given their current local knowledge. We are planning to also extend our framework so as to integrate the geometrical interpretation [22] of epistemic frames and the recent advances [35] on sheaf theoretic and homological aspects of task solvability.
@@ Appendix A Taxonomy of look-compute-move
There are several possible choices regarding the capacities of robots. The following Table 1 summarizes the most relevant variations found for robots under the look-compute-move abstraction of Definition 2.6. They are chosen in order to make a model mode or less powerful, with additional assumptions on their communication, observation and movement capacities.

### v2

@@ 4.3 Application to Surveillance
4.4 Application to Approximate Gathering
In the approximate gathering task, robots needs to ensure that their position and all other robots positions’ are eventually contained in a specific open set, called the rendezvous region. For this epistemic modeling we need to extend the language of Definition˜4.2 with additional atoms:
  • Positional atoms $pos_{\mathbf{r}}(x)\in\text{P}_{g(X)}\subseteq X$ (i.e., a finite subset of the domain of the observation function $g$ in Definition˜2.1) with $|\text{P}_{g(X)}|\in\mbox{$\mathbb{N}$}$;
  • Initial position atoms $init(pos_{\mathbf{r}}(x))\in\text{P}_{g(X)}$, meaning the starting position of robot $\mathbf{r}$;
  • Inclusion atoms $(x\in U)\in\text{P}_{\tau}$ representing the inclusion of points in open sets in the exploration space $\in_{x,U}$;
We define the set of Rendezvous region $\text{P}_{rz(X)}\in 2^{X}$, which are a a finite set of regions of space such that $\forall U,V\in\text{P}_{rz(X)}:U\cap V=\varnothing$ and $|\text{P}_{rz(X)}|\in\mbox{$\mathbb{N}$}$. $\vec{x}\in\text{P}_{X}^{\Pi}$ denotes a vector with a position for each robot. The following grammar defines temporal epistemic logic formulas for approximate gathering.
@@ 4.4 Application to Approximate Gathering
where $p\in\text{P}_{\tau}\cup\text{P}_{g(X)}$ , $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$. When obvious from the context, we will omit the subscript $X$.
The semantics of $\mathcal{L}_{gather}$ is analogous to the one of $\mathcal{L}_{exp}$ in Definition˜4.3, except for the newly introduced atoms:
  • $\mathcal{I},(\sigma,[t])\models pos_{\mathbf{r}}(x)\text{ iff }pos_{\mathbf{r}}(x)\in\pi(\sigma,[t])(pos_{\mathbf{r}}(x))$
  • $\mathcal{I},(\sigma,[t])\models init(pos_{\mathbf{r}}(x))\text{ iff }pos_{\mathbf{r}}(x)\in\pi(\sigma,0)(pos_{\mathbf{r}}(x))$
[…]
Positional atoms $pos_{\mathbf{r}}(x)$ mean that robot $\mathbf{r}$ is at actual position $\mathbf{x}$ discretized to $x$. As $g$ is surjective, it is possible that $g_{\mathbf{r}}(\mathbf{x})=g_{\mathbf{r}}(\mathbf{x}^{\prime})$ even though $\mathbf{x}\neq\mathbf{x}^{\prime}$, meaning two different actual robot positions result in the same position for that robot. We assume that robots know their own position: $pos_{\mathbf{r}}(x)\leftrightarrow K_{\mathbf{r}}pos_{\mathbf{r}}(x)$.
The approximate gathering is solved if all robots’ positions are contained in the same rendezvous region forever, assuming each robot starts in one rendezvous region. To do so, they need to agree on one rendezvous region and move here. They are allowed to change the rendezvous region finitely often, but eventually need to stick to one forever. We require it to be impossible for any robot to intersect two rendezvous regions at the same time, i.e., $pos_{r}(x)\wedge(x\in U)\implies(\bigwedge_{U\neq V\in\text{P}_{rz(X)}}\neg(x\in V))$.
Definition 4.34 (Approximate gathering).
System $\mathcal{I}$ is consistent with approximate gathering iff the following two conditions hold:
  • Eventual Gathering: There is a rendezvous region where all robots eventually meet forever
    $\mathcal{I}\models\bigvee_{U\in\text{P}_{rz(X)}}\bigwedge_{\mathbf{r}\in\Pi}\Diamond\Box\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U)$
  • Starting Validity: If all robots are already contained in a rendezvous region initially, that is also the final rendezvous region
    $\mathcal{I}\models\bigwedge_{\begin{subarray}{c}U\in\text{P}_{rz(X)},\vec{x}\in\text{P}_{g(X)}^{\Pi}\end{subarray}}\bigl{(}(\bigwedge_{r\in\Pi}init(pos_{\mathbf{r}}(x_{\mathbf{r}})))\wedge(x_{\mathbf{r}}\in U)\bigr{)}\rightarrow\bigl{(}\Diamond\Box\bigvee_{\vec{y}\in\text{P}_{g(X)}^{\Pi}}(\bigwedge_{\mathbf{r}\in\Pi}pos_{\mathbf{r}}(y_{\mathbf{r}}))\wedge(y_{\mathbf{r}}\in U)\bigr{)}.$
These conditions strongly resemble the stabilizing agreement conditions, epistemically characterized in [8]. It is defined by the following formulas, where $\mathcal{V}$ is the set of values to choose from, the atom $choose_{a}(v)$ denotes that agent $a$ chooses value $v$. $decide_{a}(v)$ is defined as $\Box choose_{a}(v)$, where $a,b$ are agents, i.e., robots with no dynamical constraints:
@@
System $\mathcal{I}$ is consistent with stabilizing agreement if (Agreement) and (Validity) hold.
  • Agreement: There is a value such that every agent decides on that value
    $\mathcal{I}\models\bigvee_{v\in\mathcal{V}}\bigwedge_{a\in\Pi}\Diamond decide_{a}(v)$
  • Validity: An agent can only choose a known initial value of some agent
    $\mathcal{I}\models\bigwedge_{v\in\mathcal{V}}\bigl{(}choose_{a}(v)\rightarrow K_{a}\bigvee_{b\in\Pi}init_{b}(v)\bigr{)}$
Observe that covering a rendezvous region can be seen as that robot choosing that region. Rendezvous regions are the objects the robots choose from. As each robot initially sits in exactly one region, that region can be seen as the problem input. The last obstacle is reducing the strong validity, as its called in the distributed computing literature (see for example [9]), to the weak validity of approximate gathering.
@@
Consider a model $\mathcal{I}$ with a space $X$, together with the language defined Definition˜4.33, an arbitrary run $\sigma$ from $\mathcal{I}$. If $\mathcal{I}$ is consistent with stabilizing agreement, then it is also consistent with approximate gathering.
Consider setting the set of possible choices $\mathcal{V}$ to be the set of possible rendezvous regions $\text{P}_{rz(X)}$, setting $choose_{\mathbf{r}}(U)\leftrightarrow\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U)$ and $decide_{\mathbf{r}}(U)=\Box choose_{\mathbf{r}}(U)$, and setting $init_{\mathbf{r}}(U)\leftrightarrow\bigvee_{x\in\text{P}_{g(X)}}init(pos_{\mathbf{r}}(x))\wedge(x\in U)$.
Proof 4.37.
We start by substituting for choosechoose and initinit as defined in Definition˜4.35. For (Agreement) we immediately arrive at (Eventual Gathering):
$\mathcal{I}\models\bigvee_{U\in\text{P}_{rz(X)}}\bigwedge_{\mathbf{r}\in\Pi}\Diamond\Box\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U).$ (11)
For (Starting Validity) we need to provide more reasoning, consider the following, where we already prepended the validity with a conjunction over all robots:
$\begin{split}\mathcal{I}\models\bigwedge_{\begin{subarray}{c}\mathbf{r}\in\Pi,\\ U\in\text{P}_{rz(X)}\end{subarray}}\bigl{(}(&\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in U))\rightarrow\\ (K_{\mathbf{r}}&\bigvee_{\mathbf{r}^{\prime}\in\Pi}\bigvee_{x\in\text{P}_{g(X)}}init(pos_{\mathbf{r}^{\prime}}(x))\wedge(x\in U))\bigr{)}.\end{split}$ (12)
@@
for any $U\in\text{P}_{rz(X)}$ and any $\vec{x}\in\text{P}_{g(X)}^{\Pi}$. Choose any $U$, $\vec{x}$ which satisfies Equation˜13, meaning a starting position vector that intersects over $U$. Note that there is only one such $U$ as initial rendezvous regions are unique.
Observe that, as robots know their own initial positions, Equation˜13 implies the antecedent in Equation˜12 for all robots $\mathbf{r}$:
@@
Therefore the consequent in Equation˜12 is true for only one choice of $U$ (the same that satisfies Equation˜13) and false for all other $V\in\text{P}_{rz(X)}$.
By Equation˜11 there exists some $V$, $t$ such that
$\mathcal{I},(\sigma,t)\models\bigvee_{x\in\text{P}_{g(X)}}pos_{\mathbf{r}}(x)\wedge(x\in V)$ (16)
holds for all $t^{\prime}>t$ and for all robots $r$. We see that Equation˜16 is precisely the antecedent of Equation˜12, and, as Equation˜12 is a validity, it is satisfied by the fixed choice for $V$, $t$. Since the consequent in Equation˜12 is true only for $U$ as derived in Equation˜15 (but Equation˜12 is a validity), Equation˜16 has to be false for any $W\neq U$. But by the previous reasoning we know that Equation˜16 is true for at least one $V$ for all $t^{\prime}>t$, implying that $U=V$ for all $t^{\prime}>t$. We can now derive Equation˜14 from Equation˜16:
@@
meaning the robots gather at the initially intersected rendezvous region forever.
@@ 5 Conclusions and future work
In this paper, we developed a framework that seamlessly integrates control theory, distributed systems and temporal epistemic logic in the context of multi-robot systems, with applications to exploration, gathering and surveillance tasks. We present, at the same time, new complementary perspectives for modeling distributed robot systems as dynamical systems and as state machines. The novel separation of a robot state into ontic and epistemic parts is exploited by developing a common understanding of scheduler that ties their local executions to a global time. We extend this first bridge between control theory and distributed computing by also providing access to the epistemic reasoning present in distributed computing, now adapted for multi-robot tasks and their particularities, namely the fact that robots exist in some (physical) space. This reasoning arena takes the form of an epistemic model capable of expressing the knowledge of the system at their computation operations. A temporal epistemic logic of space is employed and its usefulness made evident with applications to concrete exploration and gathering tasks, for both of which sufficient epistemic conditions are derived.
This work has multiple implications and holds the promise of future developments in exchanging tools and results across fields. One of the most striking implications is that the classical luminous robot model [10] with full visibility is very strong in a distributed computing sense, as confirmed by the existence of a whole research field on establishing and maintaining communication in multi robot systems. Some, such as [6, 7] have leveraged epistemic planning and action models in order to keep a live communication. As we have identified a sufficient condition for approximate gathering, we can directly judge the correctness of an action model, and therefore the correctness of a protocol, by its ability to reach a given specification in epistemic terms.
The results of this paper have already been extended for the task of surveillance, as can be found in Section˜4.3, where a group of robots must monitor a region of space to be absent of intruders. Other interesting extensions include (but are not limited to) considering other robot tasks and their similarities to distributed tasks, as well as extending the logic framework with deontic modalities expressing robots’ permissions and obligations during protocol execution given their current local knowledge. Furthermore, our framework is capable of integrating the geometrical interpretation [21] of epistemic frames and the recent advances [34] on sheaf theoretic and homological aspects of task solvability, however this is out of scope for this paper, and could be the focal interest of follow-up work.
@@ Appendix A Taxonomy of look-compute-move
There are several possible choices regarding the capacities of robots. The following Table˜1 summarizes the most relevant variations for robots under the look-compute-move abstraction. They are chosen in order to make a model mode or less powerful, with additional assumptions on their communication, observation and movement capacities.